HelpMePlz:

use the information given for the year 1995 and 2000, letting x=5 represent 1995, x=10 represent 2000, and y the number of post offices, to write an equation that models the data. Graph (amount of post offices)- 1995 has 28,392, 2000 has 27,876

Champion:

There are many ways to model a set of data, linear, exponential, square, sqaure-root. With only two points anything goes! If it is linear, then use y=ax+b If it is exponential, try y=ab^x If it is square, try y=ax^2+bx+c (but you have to assume one of a, b or c) and so on. Find the unknown constants using the given data and you'll have it!